A survey on British Social Attitudes asked respondents if they had ever boycotted goods for ethical reasons (Statesman, January 28, 2008). The survey found that 23% of the respondents have boycotted goods for ethical reasons.

In a sample of ten British citizens, what is the expected number of people that have boycotted goods for ethical reasons? Also find the standard deviation.
Answer #1

p=0.23

q=1-p =1-0.23. =0.77

n=10

1)Expected number of people that have boycotted goods for ethical reasons

=np

=0.23 * 10

=2.3

=3

2) standard deviation =√(npq)

=√(10* 0.23 * 0.77)

=1.771